OpenStudy (anonymous):

p[(5/p)-0.004]=-1 solve for p

jimthompson5910 (jim_thompson5910):

Start by distributing p[(5/p)-0.004]=-1 p*(5/p)-0.004p=-1 5p/p-0.004p=-1 1-0.004p=-1 Can you take it from here?

OpenStudy (anonymous):

-2/-0.004=p?

OpenStudy (anonymous):

500?

jimthompson5910 (jim_thompson5910):

oh wait, i made a mistake (my bad)

jimthompson5910 (jim_thompson5910):

p[(5/p)-0.004]=-1 p*(5/p)-0.004p=-1 5p/p-0.004p=-1 5-0.004p=-1

jimthompson5910 (jim_thompson5910):

so it should be 5-0.004p=-1

OpenStudy (anonymous):

-6/-0.004=1500=p?

jimthompson5910 (jim_thompson5910):

yes, p = 1500

jimthompson5910 (jim_thompson5910):

nice job

OpenStudy (anonymous):

thankyou!!

jimthompson5910 (jim_thompson5910):

np